A ball is drawn at random from a box containing 10 red balls, 30 white balls, 20 blue balls and 15 black balls.Obtain the probability that the ball drawn is:
(1)black or red
(2) not red or blue
(3)not blue
(4) white
(5) red, white,blue
1
Expert's answer
2020-05-01T18:14:29-0400

The total number of ball in the box , "n(s)=10+30+20+15=75" .

Let, red ball denoted by "R"

White ball denoted by "W"

Blue ball denoted by "B"

Black ball denoted by "Bl" .


1) "P(R \ or Bl)=P(B)+P(Bl)"


"=\frac{10}{75}+\frac{15}{75}=\frac{25}{75}=\frac{1}{3}"


2) Let not red is denoted by "\sim R" .

"P(\sim R \ or B)=P(\sim R \ \cup B)=P(\sim R)+P(B)-P(\sim R\ \cap B)"

"=1-P(R)+P(B)-P(B)" [since "B\sub\ ( \sim R)" ]

"=1-\frac{10}{75}=\frac{65}{75}=\frac{13}{15}" .


3) "P(not \ B)=1-P(B)"

"=1-\frac{20}{75}=\frac{55}{75}=\frac{11}{15}" .


4) "P(W)=\frac{30}{75}=\frac{2}{5}"


5) "P( R \ \cup W\ \cup \ B)=P(R)+P(W)+P(B)"

(As R,W,B are disjoint)

"=\frac{10}{75}+\frac{30}{75}+\frac{20}{75}=\frac{60}{75}=\frac{4}{5}" .